Write the following in logarithmic form: 2e^x=1
Please Help with this equation

Found 2 solutions by Alan3354, FrankM:
Answer by Alan3354(69443)

Write the following in logarithmic form: 2e^x=1
----------------
Divide by 2.
---
The log is the exponent.
e is the base.

Answer by FrankM(1040)

2e^x=1
e^x=1/2
x = ln 1/2